OpenStudy (anonymous):

Show how to find the derivative of f(t)= (t^2+1)/t using the quotient rule. Then show how to find the derivative of the same function without using the quotient rule and first changing the form of the function. I get two different answers when I do this! HELP MEEEEEEEE

OpenStudy (amistre64):

1/t = t^-1

myininaya (myininaya):

\[f'(t)=\frac{(t^2+1)'t-(t^2+1)(t)'}{t^2}\]

OpenStudy (amistre64):

\[[t+t^{-1}]'=1-t^{-2}\]

myininaya (myininaya):

\[=\frac{(2t+0)t-(t^2+1)(1)}{t^2}\]

myininaya (myininaya):

\[=\frac{2t^2-(t^2+1)}{t^2}=\frac{2t^2-t^2-1}{t^2}=\frac{t^2-1}{t^2}=\frac{t^2}{t^2}-\frac{1}{t^2}\] same as what amistre has

OpenStudy (amistre64):

\[[t+t^{-1}]'=1-t^{-2}\] \[[t+t^{-1}]'=\frac{t^2-1}{t^{2}}\]

myininaya (myininaya):

1-t^{-2}

myininaya (myininaya):

or he can rewrite his to match mine

OpenStudy (amistre64):

lol yay!!

OpenStudy (anonymous):

this a repeat?

OpenStudy (amistre64):

its all a repeat ;)

OpenStudy (anonymous):

No, it's not a repeat, on the first one, I didn't put the numerator in parenthesis, so it was done wrong.

myininaya (myininaya):

no they are talking about the question

myininaya (myininaya):

they seen it before and i think i seen it too

OpenStudy (anonymous):

I don't understand how to do it without the quotient rule though! Can someone explain that!

OpenStudy (anonymous):

@nls11, it is a good thing to remember (memorize if you will) three fact: \[\frac{d}{dx}\frac{1}{x}=-\frac{1}{x^2}\] \[\frac{d}{dx}\sqrt{x}=\frac{1}{2\sqrt{x}}\]and their chain rule cousins \[\frac{d}{dx}\frac{1}{f(x)}=-\frac{f'(x)}{f^2(x)}\] and \[\frac{d}{dx}\sqrt{f(x)}=\frac{f'(x)}{2\sqrt{f(x)}}\]

myininaya (myininaya):

nls, that what amistre was doing above he wrote as two fractions

OpenStudy (anonymous):

math teachers love putting these on exams, and it saves you from rewriting everything using stupid exponents and then using the power rule and getting rid of the exponents

myininaya (myininaya):

\[\frac{t^2+1}{t}=\frac{t^2}{t}+\frac{1}{t}\]

myininaya (myininaya):

\[=t+\frac{1}{t}\] then we find the derivative of that since this is what your question asks